Shift summary for evening 7/11/99

Hogan Nguyen (hogann@fnal.gov)
Mon, 12 Jul 1999 00:21:35 -0500

Shift Summary for evening shift, 07/11/99

On Shift: Pat Toale, Hogan Nguyen

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 8.0
Hours of E832 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: DA problem

========================================================================
Runs ended this shift:

Run 13857 - BEAM_832 (24.1 M events)
Run 13858 - BEAM_832 (very short run)
Run 13859 - BEAM_832 (very short run)

Tapes brought to FCC this shift: QK*116-118, QKB040
========================================================================

Accesses during shift: none
========================================================================

Overview of shift: 832 data taking at ~6.5 E12

Ran into problems with monitor and filter CPUs taking
too long to process.

Occasional ktevana errors, HCC inefficiency warnings,
Pipeline resetting requests.

Missing wire in DC4Y Kumquat latch illumination

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13857 16:07 Tapes are full. This run stops. We load a new
tape set.

13858 16:26 Start beam832 run 13858

13858 17:20 DDD3 veto is asserted frequently, thus inhibitting
triggers and resulting in very bad livetime.
Called Ed. We traced this eventually to a monitor
process taking twice as long as it's supposed to.

Stopped this run. Ed increases the prescale for
trigger 2 events going into the monitor plane.

13859 18:43 Start new run. For this run, we "brute force"
initialized the L3 processes. The monitor processes
are now O.K., but there is now a filter process
taking too much time.

After consulting with Ed, we stop this run and
reinitialized with drt_fresh_start

13860 19:08 Start new beam832 run. Filter and monitoring processes
are o.k.

13860 20:30-21:03 No beam due to Tevatron abort.

13860 23:00 We notice that for this entire run, wire 43 is
dead in the kumquat illumination. This is used in
the YTF. Reloading the kumquats did not help.

Called Rick Kessler. The e832#kqke3 efficiency is
still good since the trigger allows 1 missing hit in
the chamber. We decide to leave it alone until the
Tuesday shutdown.

The behaviour of this wire is intermittent. It is
mostly dead in runs 13847,58,59,60.

========================================================================
Follow up Notes on Previous Shift

Keep an eye on the +/- 24V supply to RR160.

========================================================================
Items to Follow up on Next Shift

Keep an eye on DC4Y kumquat illumination.